OpenStudy (anonymous):

0.5x+6 is greater than or equal to 0.9x-9

OpenStudy (anonymous):

0.5x+6>=0.9x-9 0.5x-0.9x>=-9-6 x(0.5-0.9)>=-15 x>=-15/(0.5-0.9) x>=37.5